Z1 and Z2 are first connected in series then in parallel Express the combined impendances in each casein form of x +jy ; if Z1 = 4-j3 and Z2= 5+j2


1
Expert's answer
2021-04-14T07:43:39-0400

In case of series connection:


"Z=Z_1+Z_2=4-j3+5+j2=(9-j)\Omega."

For parallel connection:


"Z=\frac{Z_1Z_2}{Z_1+Z_2}=\frac{(4-j3)(5+j2)}{4-j3+5+j2}=(2.94-j0.451)\Omega."